OpenStudy (masterdestroyer55):

What is the surface area of a rectangular prism with a length of 3 mm, a width of 5 mm, and a height of 8 mm? ________mm2 158?

OpenStudy (masterdestroyer55):

@imqwerty @dayakar @mathstudent55

OpenStudy (masterdestroyer55):

@Qwertty123

OpenStudy (masterdestroyer55):

@King.Void.

OpenStudy (ltrout):

Do you know how to find surface area?

OpenStudy (masterdestroyer55):

yes. add them all up then divide by 3?

OpenStudy (ltrout):

The formula is \[A = 2(wl + hl + hw)\]

OpenStudy (masterdestroyer55):

so then A = 158?

OpenStudy (ltrout):

One sec let me get a calculator lol

OpenStudy (masterdestroyer55):

A=2(wl+hl+hw)=2·(5·3+8·3+8·5)=158 is it correct?

OpenStudy (ltrout):

Yes that's correct :)

OpenStudy (ltrout):

Sorry still too early for me to be doing math in my head

OpenStudy (masterdestroyer55):

ok i got confused, because the last question i had, the answer was also 158. i just wanted to make sure

OpenStudy (masterdestroyer55):

i have more, but im going to open a new question first.

OpenStudy (mathstudent55):

Correct. Another way: Lateral area = perimeter of the base * height Base area = 2 * length * width of the base LA = (3 + 3 + 5 + 5)mm * 8 mm = 128 mm^2 BA = 2 * 3 mm * 5 mm = 30 mm^2 Total surface area = LA + BA = 128 mm^2 + 30 mm^2 = 158 mm^2

OpenStudy (ltrout):

Which would you say is quicker @mathstudent55

OpenStudy (mathstudent55):

I explained each step which makes it longer. Here it is in a quicker one-line expression: SA = Ph + 2lw = (3 + 3 + 5 + 5)8 + 2 * 3 * 5 = 158 I think that your method is still quicker, but there are occasions in which you are asked for just the lateral area, so it is also good to know how to find the lateral area.